OpenStudy (anonymous):

The perimeter of a rectangle is 24ft. The length is 4ft longer than the width. Find the dimensions. write a system linear equation. x is the length and y is the width. Help

hartnn (hartnn):

since x is 4 feet longer than width,u can write x-y=4 perimeter=2(x+y)=24 to give x+y=12 can u find x and y from these 2 equations?

OpenStudy (anonymous):

x - y = 4 x = 4 + y So you get x in terms of y now. Plug in this value of x in x + y = 12 and get the answer.

OpenStudy (anonymous):

Yes. Thank you